# Michael–Simon inequality for anisotropic energies close to the area via multilinear Kakeya-type bounds

### v1

@@ 3.3. Complementary inequalities
With $G:=G^{x}\cap G^{y}$, we deduce that
$\int_{\mathbb{R}^{2}}\chi\min\{S^{P},T^{P}\}$ $\leq\int_{G}\chi\min\{S_{f}^{P},T_{f}^{P}\}+\int_{\mathbb{R}^{2}\setminus G}\min\{S_{f}^{P},T_{f}^{P}\}+\int_{\mathbb{R}^{2}}(S_{d}^{P}+T_{d}^{P})$
$\leq\int_{G}\chi\sqrt{S_{f}^{P}T_{f}^{P}}+\int_{\mathbb{R}^{2}\setminus G^{x}}S_{f}^{P}+\int_{\mathbb{R}^{2}\setminus G^{y}}T_{f}^{P}+\int_{\mathbb{R}^{2}}(S_{d}^{P}+T_{d}^{P})$
@@
$\leq\int_{G}\chi\sqrt{S_{f}^{P}T_{f}^{P}}+\Lambda\int_{\mathbb{R}^{2}}(\lvert\operatorname{div}S\rvert+\lvert\operatorname{div}T\rvert)+\int_{\mathbb{R}^{2}}(S^{N}+T^{N}),$
where we used Eq. 20 and its analogue for $T$ in the last inequality.
Finally, we have $S_{f}^{P}T_{f}^{P}\leq C\det(S_{f},T_{f})$ on $G$ by the elementary bound
$(v^{x}-|v^{y}|)(w^{y}-|w^{x}|)\leq C\det(v,w)$
for vectors $v\in\overline{C^{x}}$ and $w\in\overline{C^{y}}$ (note that $S_{f}\in\overline{C^{x}}$ on $G$, except possibly on the negligible set $\mathbb{Z}\times\mathbb{R}$, and similarly $T_{f}\in\overline{C^{y}}$). In turn, this bound follows from the fact that, assuming $|v|=|w|=1$, $\det(v,w)$ is comparable with $\min\{|v-w|,|v+w|\}$, while $v^{x}-|v^{y}|$ is comparable with $\operatorname{dist}(v,\partial C^{x})$, and similarly $w^{x}-|w^{y}|$ is comparable with $\operatorname{dist}(w,\partial C^{y})$.
From this remark and Cauchy–Schwarz it follows that
$\int_{G}\chi\sqrt{S_{f}^{P}T_{f}^{P}}\leq C\Big(\int_{\mathbb{R}^{2}}\chi^{2}\Big)^{1/2}\Big(\int_{G}\det(S_{f},T_{f})\Big)^{1/2}.$
However, since $G=\bigcup_{j,k}G_{jk}$, by Eq. 16 we have
[…]
This concludes the proof of 3.1.
@@ 4. Proof of Michael–Simon for anisotropies close to the area
is symmetric and positive semidefinite. Note carefully that the same matrix $\mathcal{A}$ given by Eq. 22Eq. 24 is not symmetric for a general $F$, nor it has (formally) nonnegative determinant in general. However, we now show that one can always reduce to the case where the diagonal entries are nonnegative, assuming that $F$ is convex. In fact, the following holds in arbitrary dimension (when k=n1k=n-1), with the same proof.
@@
when $F$ is the area, which implies that
@@
Finally, we set $E:=\pi_{x,y}(\{\theta>0\})$, which is an analytic set; as such, we can find a Borel set $E^{\prime}\supseteq E$ such that $\mathcal{L}^{2}(E^{\prime}\setminus E)=0$. Observing that $S,T$ are concentrated on $E^{\prime}$ and applying 3.1 with $\chi:=\bm{1}_{E}$, we get
@@
Recalling that $(\nu^{x})^{2}+(\nu^{y})^{2}=1-(\nu^{z})^{2}$ and Eq. 27, the fact that $\gamma>4$ implies that $(\nu^{z})^{2}-(\frac{1}{4}+\frac{1}{\gamma})(1-(\nu^{z})^{2})$ is at least $\frac{1}{3}-(\frac{1}{4}+\frac{1}{\gamma})\frac{2}{3}=:2c>0$ in average, and hence
$c|V|(\mathbb{R}^{3})\leq C\mathcal{L}^{2}(E)^{1/2}|V|(\mathbb{R}^{3})+(C\mathcal{L}^{2}(E)^{1/2}+1)|\delta^{F}V|(\mathbb{R}^{3}),$
provided that $\varepsilon<c$. However, since $\mathcal{L}^{2}(E)\leq\mathcal{H}^{2}(\{\theta>0\})=\eta_{0}$, we can now choose $\eta_{0}$ such that $C\sqrt{\eta_{0}}\leq\frac{c}{2}$ and deduce that $|V|(\mathbb{R}^{3})\leq C|\delta^{F}V|(\mathbb{R}^{3})$ (for a possibly different $C$). This completes the proof of 1.3.
@@
$\nu^{x}\partial_{x}F(\nu)\geq 0,\quad\nu^{y}\partial_{y}F(\nu)\geq 0,\quad\nu^{z}\partial_{z}F(\nu)\geq 0,$
@@
Taking $\mathcal{A}$ as in the proof of 1.3, we claim that $\mathcal{A}\in\mathcal{M}^{+}$. Indeed, let us disintegrate $V$ with respect to the projection $\Pi$. Writing
$V(x,y,z,\nu)=\Pi_{*}V(x,y)\otimes\lambda_{x,y}(z,\nu)$
for a family of probability measures $\lambda_{x,y}$ on $\mathbb{R}\times\mathbb{S}^{2}$, we have $\mathcal{A}=\Lambda\Pi_{*}V$, where
$\Lambda(x,y)=\int_{\mathbb{R}\times\mathbb{S}^{2}}\begin{pmatrix}\nu^{y}\partial_{y}F(\nu)+\nu^{z}\partial_{z}F(\nu)&-\nu^{x}\partial_{y}F(\nu)\\ -\nu^{y}\partial_{x}F(\nu)&\nu^{x}\partial_{x}F(\nu)+\nu^{z}\partial_{z}F(\nu)\end{pmatrix}\,d\lambda_{x,y}(z,\nu).$
This gives immediately $\mathcal{A}_{x}^{x},\mathcal{A}_{y}^{y}\geq 0$. Finally, we have $\det(\Lambda)\geq 0$ pointwise since
$\det(\Lambda(x,y))$ $)\displaystyle=\iint_{(\mathbb{R}\times\mathbb{S}^{2})^{2}}[(\nu^{y}\partial_{y}F(\nu)+\nu^{z}\partial_{z}F(\nu))(\tilde{\nu}^{x}\partial_{x}F(\tilde{\nu})+\tilde{\nu}^{z}\partial_{z}F(\tilde{\nu}$))
@@
thanks to Eq. 30. Hence, we obtain
$\sqrt{\det(\Lambda(x,y))}\geq\int_{(\mathbb{R}\times\mathbb{S}^{2})^{2}}\nu^{z}\partial_{z}F(\nu)\,d\lambda_{x,y}(z,\nu),$
giving
$\sqrt{\det(\mathcal{A})}(\mathbb{R}^{3})=\int_{\mathbb{R}^{2}}\sqrt{\det(\Lambda)}\,d\Pi_{*}V\geq\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\nu^{z}\partial_{z}F(\nu)\,dV(p,\nu).$
We now conclude as in the proof of 1.3, using Corollary 3.6 in place of 3.1.

### v2

@@ 3.3. Complementary inequalities
Letting
$\mathcal{R}^{x}:=\bigcup_{j,\ell}\mathcal{R}_{j,\ell}^{x},\quad\mathcal{R}^{y}:=\bigcup_{k,m}\mathcal{R}_{k,m}^{y},$
and $\mathcal{R}:=\mathcal{R}^{x}\cap\mathcal{R}^{y}$, we deduce that
$\int_{\mathcal{R}}\chi\min\{S^{P},T^{P}\}$ $\leq\int_{G}\chi\min\{S_{f}^{P},T_{f}^{P}\}+\int_{\mathcal{R}\setminus G}\min\{S_{f}^{P},T_{f}^{P}\}+\int_{\mathcal{R}}(S_{d}^{P}+T_{d}^{P})$
$\leq\int_{G}\chi\sqrt{S_{f}^{P}T_{f}^{P}}+\int_{\mathcal{R}^{x}\setminus G^{x}}S_{f}^{P}+\int_{\mathcal{R}^{y}\setminus G^{y}}T_{f}^{P}+\int_{\mathcal{R}}(S_{d}^{P}+T_{d}^{P}).$
Also, by Proposition 3.10 we have
$\int_{\mathcal{R}^{x}\setminus G^{x}}S_{f}^{P}\leq\int_{\mathcal{R}^{x}}[S_{f}^{N}+\lvert\operatorname{div}S\rvert$],
while Proposition 3.13 gives
$\int_{\mathcal{R}^{x}}S_{d}^{P}\leq\int_{\mathbb{R}^{2}}[C^{\prime}S_{d}^{N}+C\lvert\operatorname{div}S\rvert$],
and similarly for $T$. Once we sum these bounds, we get
$\int_{\mathcal{R}}\chi\min\{S^{P},T^{P}\}$ $\leq C\|\chi\|_{L^{2}}\Big[\int_{\mathbb{R}^{2}}(|S|+\lvert\operatorname{div}S\rvert)\Big]^{1/2}\Big[\int_{\mathbb{R}^{2}}(|T|+\lvert\operatorname{div}T\rvert)\Big]^{1/2}$
$\quad+C\int_{\mathcal{R}}(\lvert\operatorname{div}S\rvert+\lvert\operatorname{div}T\rvert)+C^{\prime}\int_{\mathcal{R}}(S^{N}+T^{N}),$
where we can take $C^{\prime}=1$ if $S^{x},T^{y}\geq 0$. Thus, we are left to bound the integral
$\int_{\mathbb{R}^{2}\setminus\mathcal{R}}\min\{S^{P},T^{P}\}\leq\int_{\mathbb{R}^{2}\setminus\mathcal{R}^{x}}S^{P}+\int_{\mathbb{R}^{2}\setminus\mathcal{R}^{y}}T^{P}.$
We will just bound the last integral of $S^{P}$, as the one for $T^{P}$ is analogous.
Given $j\in\mathbb{Z}$, $\delta>0$, and a smooth $0\leq\chi\leq 1$ vanishing precisely on $F$, we now let
$\mathfrak{S}:=\bm{1}_{\mathcal{U}_{j}^{x}\setminus\bigcup_{\ell}\mathcal{R}_{j,\ell}^{x}}(S+\delta\chi^{2}U).$
It is immediate to check that $\mathfrak{S}$ has summable divergence, so that we can identify it with a normal $1$-current (still denoted by $\mathfrak{S}$), with
$\mathbb{M}(\partial\mathfrak{S})=\int_{\mathbb{R}^{2}}\lvert\operatorname{div}\mathfrak{S}\rvert$.
By [15, Theorem C] and the Fleming–Rishel formula (see, e.g., [15, p. 849]), we can write
(19) $\mathfrak{S}=\int_{\Sigma}\llbracket\gamma_{\sigma}\rrbracket\,d\lambda(\sigma)+\int_{\Xi}\llbracket\gamma_{\xi}\rrbracket\,d\mu(\xi),$
where $(\gamma_{\sigma})_{\sigma\in\Sigma}$ is a family of (injective) arcs of finite length, while $(\gamma_{\xi})_{\xi\in\Xi}$ is a family of integral cycles, i.e., finite or countable unions of loops of finite total length, with respect to nonnegative measures $\lambda,\mu$ on $\Sigma,\Xi$, respectively, such that
$\mathbb{M}(\partial\mathfrak{S};\Omega)=\int_{\Sigma}\mathbb{M}(\partial\llbracket\gamma_{\sigma}\rrbracket;\Omega)\,d\lambda(\sigma)$
for any open $\Omega\subseteq\mathbb{R}^{2}$. In particular, taking $\Omega$ to be the interior of $\mathcal{U}_{j}^{x}$, we have
@@
$\lambda(\tilde{\Sigma})\leq\int_{\Omega}\lvert\operatorname{div}(S+\delta\chi^{2}U)\rvert$,
where $\tilde{\Sigma}\subseteq\Sigma$ is the subset of indices $\sigma\in\Sigma$ such that one of the endpoints of $\gamma_{\sigma}$ is in $\Omega$, since we have $\mathbb{M}(\partial\llbracket\gamma_{\sigma}\rrbracket;\Omega)\geq 1$ for each $\sigma\in\tilde{\Sigma}$ and $\lvert\operatorname{div}\mathfrak{S}\rvert\leq\lvert\operatorname{div}(S+\delta\chi^{2}U)\rvert$ on this $\Omega$. Letting
$H(\gamma):=\int_{a}^{b}[(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{+}-C^{\prime}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}]\,dt$
with $C^{\prime}$ as in 3.12, for each arc $\gamma:[a,b]\to\mathbb{R}^{2}$ (and similarly for a union of loops) we claim that
$H(\gamma_{\sigma})\leq\Lambda\bm{1}_{\tilde{\Sigma}}(\sigma),\quad H(\gamma_{\xi})\leq 0$
for any $\sigma\in\Sigma$ and $\xi\in\Xi$, which immediately gives the bound
$\int_{\mathcal{U}_{j}^{x}}(\mathfrak{S}^{x}-|\mathfrak{S}^{y}|)^{+}\leq C^{\prime}\int_{\mathcal{U}_{j}^{x}}[(\mathfrak{S}^{x}-|\mathfrak{S}^{y}|)^{-}+\Lambda\lvert\operatorname{div}(S+\delta\chi^{2}U)\rvert$],
and similarly for $T$; as usual, we can take $C^{\prime}=1$ if $S^{x},T^{y}\geq 0$. In turn, this proves 3.1 once we let $\delta\to 0$ and sum over $j$ and $k$.
To check the claim, we note that each $\gamma=\gamma_{\sigma}$ and $\gamma=\gamma_{\xi}$ is part of a curve of the types (ii)–(iv) (for a.e. $\sigma$ and a.e. $\xi$), since Eq. 19 holds without cancellation and thus $\dot{\gamma}(t)$ must be parallel to $\mathfrak{S}(\gamma(t))$ (and thus to $U(\gamma(t))$ as $\delta>0$) for a.e. $t$.
The claim follows immediately for all $\xi\in\Xi$, since part (i) of 3.12 can be applied to loops. Moreover, if $\sigma\not\in\tilde{\Sigma}$ then $\gamma_{\sigma}$ is a full curve of type (ii)–(iv), but types (iii)–(iv) are ruled out since $\gamma_{\sigma}$ has finite length and is not a loop. Hence, in this case $\gamma_{\sigma}$ is a full curve of type (ii), in which case we can apply again part (i) of 3.12.
To conclude, for an arc $\gamma_{\sigma}$ with $\sigma\in\tilde{\Sigma}$ the claim $H(\gamma_{\sigma})\leq\Lambda=\Lambda\bm{1}_{\tilde{\Sigma}}(\sigma)$ follows immediately from part (ii) of 3.12.
@@ 4. Proof of Michael–Simon for anisotropies close to the area
is symmetric and positive semidefinite. Note carefully that the same matrix $\mathcal{A}$ given by Eq. 21Eq. 23 is not symmetric for a general $F$, nor it has $\mathcal{A}_{x}^{x},\mathcal{A}_{y}^{y}\geq 0$, nor (formally) nonnegative determinant in general.
However, we now show that one can always reduce to the case where the diagonal entries are nonnegative, assuming that $F$ is convex. In fact, the following holds in arbitrary dimension (when k=n1k=n-1), in the sense that $B_{F}(\nu)$ from Eq. 8 can be assumed to have nonnegative diagonal entries, with the same proof.
@@
when $F$ is the area (i.e., $F|_{\mathbb{S}^{2}}=1$), which implies that
@@
Finally, we set $E:=\pi_{x,y}(\{\theta>0\})$, which is an analytic set; as such, we can find a Borel set $E^{\prime}\supseteq E$ such that $\mathcal{L}^{2}(E^{\prime}\setminus E)=0$. Observing that $S,T$ are concentrated on $E^{\prime}$ and applying 3.1 with $\chi:=\bm{1}_{E^{\prime}}$, we get
@@
Recalling that $(\nu^{x})^{2}+(\nu^{y})^{2}=1-(\nu^{z})^{2}$ and Eq. 26, the fact that $\gamma>4$ implies that $(\nu^{z})^{2}-(\frac{1}{4}+\frac{1}{\gamma})(1-(\nu^{z})^{2})$ is at least $\frac{1}{3}-(\frac{1}{4}+\frac{1}{\gamma})\frac{2}{3}=:4c>0$ in average, and hence
$2c|V|(\mathbb{R}^{3})\leq C\mathcal{L}^{2}(E)^{1/2}|V|(\mathbb{R}^{3})+(C\mathcal{L}^{2}(E)^{1/2}+1)|\delta^{F}V|(\mathbb{R}^{3}),$
provided that $\varepsilon<c$. However, since $\mathcal{L}^{2}(E)\leq\mathcal{H}^{2}(\{\theta>0\})=\eta_{0}$, we can now choose $\eta_{0}$ such that $C\sqrt{\eta_{0}}\leq c$ and deduce that $|V|(\mathbb{R}^{3})\leq C|\delta^{F}V|(\mathbb{R}^{3})$ (for a possibly different $C$). This completes the proof of 1.3.
@@
(29) $\nu^{x}\partial_{x}F(\nu)\geq 0,\quad\nu^{y}\partial_{y}F(\nu)\geq 0,\quad\nu^{z}\partial_{z}F(\nu)\geq 0,$
@@
Taking $\mathcal{A}$ as in the proof of 1.3 and letting $\mathcal{A}_{\varepsilon}:=\rho_{\varepsilon}*\mathcal{A}$ be a regularization by mollification, we clearly have $\mathcal{A}_{\varepsilon}\in W^{1,2}(\mathbb{R}^{2})$. Moreover, we claim that
$(\mathcal{A}_{\varepsilon})_{x}^{x}\geq 0,\quad(\mathcal{A}_{\varepsilon})_{y}^{y}\geq 0,\quad\det(\mathcal{A}_{\varepsilon})\geq 0.$
Indeed, recall that
$\mathcal{A}=\Pi_{*}[B_{F}^{\prime}(\nu)V(\cdot,\nu)],$
where
$B_{F}^{\prime}(\nu):=\begin{pmatrix}\nu^{y}\partial_{y}F(\nu)+\nu^{z}\partial_{z}F(\nu)&-\nu^{x}\partial_{y}F(\nu)\\ -\nu^{y}\partial_{x}F(\nu)&\nu^{x}\partial_{x}F(\nu)+\nu^{z}\partial_{z}F(\nu)\end{pmatrix}$,
which gives immediately $\mathcal{A}_{x}^{x},\mathcal{A}_{y}^{y}\geq 0$. In order to check that $\det(\mathcal{A}_{\varepsilon})\geq 0$ holds at any given point $(x_{0},y_{0})$, we observe that
$\mathcal{A}_{\varepsilon}(x_{0},y_{0})=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\rho_{\varepsilon}(x_{0}-x,y_{0}-y)B_{F}^{\prime}(\nu)\,dV(x,y,z,\nu).$
Thus, calling $\Lambda$ the $\nu$-marginal of $\rho_{\varepsilon}(x_{0}-x,y_{0}-y)V(x,y,z,\nu)$ (i.e., its pushforward with respect to the projection $\mathbb{R}^{3}\times\mathbb{S}^{2}\to\mathbb{S}^{2}$), we have
$\det(\mathcal{A}_{\varepsilon}(x_{0},y_{0}))$ $)\displaystyle=\iint_{(\mathbb{S}^{2})^{2}}[(\nu^{y}\partial_{y}F(\nu)+\nu^{z}\partial_{z}F(\nu))(\tilde{\nu}^{x}\partial_{x}F(\tilde{\nu})+\tilde{\nu}^{z}\partial_{z}F(\tilde{\nu}$))
@@
thanks to Eq. 29 and Eq. 30. In particular, by Eq. 29 again, we see that $\det(\mathcal{A}_{\varepsilon})\geq 0$.
3.5 now gives
$\int_{\mathbb{R}^{2}}\det(\mathcal{A}_{\varepsilon})\leq\frac{|\delta^{F}V|(\mathbb{R}^{3})^{2}}{4},$
since the divergence of each row of $\mathcal{A}_{\varepsilon}$ is bounded in $L^{1}(\mathbb{R}^{2})$ by $|\delta^{F}V|(\mathbb{R}^{3})$ (as this holds for $\mathcal{A}$). Hence, fixing $\chi:\mathbb{R}^{2}\to[0,1]$ continuous and applying Cauchy–Schwarz, we get
$\int_{\mathbb{R}^{2}}\chi\sqrt{\det(\mathcal{A}_{\varepsilon})}\leq\|\chi\|_{L^{2}}|\delta^{F}V|(\mathbb{R}^{3}).$
Next, given any measure $M$ with values in $\mathbb{R}^{2\times 2}$, writing $M=A|M|$ for some unit-valued Borel $A:\mathbb{R}^{2}\to\mathbb{R}^{2\times 2}$ we let
$\sqrt{\det(M)^{+}}:=\sqrt{\det(A)^{+}}|M|,$
which is a well-defined nonnegative measure. Since we have the tight convergence $|\mathcal{A}_{\varepsilon}|\rightharpoonup|\mathcal{A}|$, by Reshetnyak’s continuity principle we also have
$\sqrt{\det(\mathcal{A}_{\varepsilon})^{+}}\rightharpoonup\sqrt{\det(\mathcal{A})^{+}}$
as measures, since the assignment $A\mapsto\sqrt{\det(A)^{+}}$ is continuous and $1$-homogeneous on $\mathbb{R}^{2\times 2}$. Hence, we can pass to the limit and deduce that
$\int_{\mathbb{R}^{2}}\chi\,d\sqrt{\det(\mathcal{A})^{+}}\leq\|\chi\|_{L^{2}}|\delta^{F}V|(\mathbb{R}^{3}).$
By approximation, this then also holds for any Borel $\chi:\mathbb{R}^{2}\to[0,1]$. As in the proof of 1.3, we can let $\chi$ be the indicator function of $E:=\pi_{x,y}(\{\theta>0\})$. The same computation used above shows that
$\int_{\mathbb{R}^{2}}\bm{1}_{E}\,d\sqrt{\det(\mathcal{A})^{+}}=\sqrt{\det(\mathcal{A})^{+}}(\mathbb{R}^{3})\geq\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\nu^{z}\partial_{z}F(\nu)\,dV(p,\nu).$
Since the latter is $\geq c|V|(\mathbb{R}^{3})$ by assumption, we arrive at
$c|V|(\mathbb{R}^{3})\leq\mathcal{H}^{2}(\{\theta>0\})^{1/2}|\delta^{F}V|(\mathbb{R}^{3}),$
as desired.